Please help me solve this:
"A is 2 years older than B. 5 years from now, A will be 16/15 as old as B. How old is B now?"

Answer by rothauserc(4718)

we have the following equations
A = B +2
A + 5 = (16/15) * (B + 5)
A = (16B/15 + 16/3) - 5
A = 16B/ 15 + 80/15 - 75/15
A = (16B +5) / 15
15A = 16B +5
15A - 16B = 5
substitute for A in above equation
15(B+2) -16B = 5
15B + 30 -16B = 5
-B = -25
B = 25
A = 27
check answer
27 + 5 = (16/15) * (25 + 5)
32 = (16/15) * 30
32 = 32
our answer checks